OpenStudy (anonymous):

verify sin^4(x)-1 / cos^2(x) = cos^2(x)-2

OpenStudy (anonymous):

I don't understand how to get rid of the ^4

hero (hero):

Is it \[\frac{\sin^4x - 1}{\cos^2x} = \cos^2x - 2\]

OpenStudy (anonymous):

yes

hero (hero):

Okay so according to the difference of squares rule \(a^2 - b^2 = (a + b)(a - b)\) \[\frac{\sin^4x - 1}{\cos^2x} = \frac{(\sin^2x + 1)(\sin^2x - 1)}{\cos^2x}\] Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right

hero (hero):

Correction: By Pythagorean Identity we can write \[\frac{(\sin^2x + 1)(\sin^2x - 1)}{\cos^2x} = \frac{(\sin^2x + 1)(\sin^2x - 1)}{1 -\sin^2x }\] Right?

OpenStudy (anonymous):

okay yes

OpenStudy (anonymous):

i agree

hero (hero):

Actually, if we take out the negative from the denominator, we will get: \[\frac{(\sin^2x + 1)(\sin^2x - 1)}{1 -\sin^2x } = -\frac{(\sin^2x + 1)(\sin^2x - 1)}{\sin^2x - 1 }\] Okay?

OpenStudy (anonymous):

okay

hero (hero):

Which means \(\sin^2x - 1\) cancels to get: \[-(\sin^2x + 1)\]

hero (hero):

Now of course \(\sin^2x = 1 - \cos^2x\) so \(-(\sin^2x + 1) = -((1- \cos^2x) + 1)\) Do you agree?

OpenStudy (anonymous):

yes

hero (hero):

And we can remove the inner set of parentheses to simplify it to: \(-((1- \cos^2x) + 1) = -(1 - \cos^2x + 1)\) Do you see how we get the RHS from there?

OpenStudy (anonymous):

yes

hero (hero):

Okay great :) Any questions?

OpenStudy (anonymous):

nope! thank you SO much